# Negative Dependence as a toolbox for machine learning : review and new developments

### v1

@@ 11.1 The formalism of quantum physics
For simplicity, we henceforth restrict to the $2^{N}$-dimensional example $\mathbb{H}=(\mathbb{C}^{2})^{\otimes N}$, for some $N\geq 0$. This is the Hilbert space that is used in quantum computing to describe a system of $N$ qubits, i.e. $N$ particles that can each assume two states. In this finite-dimensional setting, an observable $A$ is simply a self-adjoint matrix, so that it diagonalizes in an orthonormal basis $A=\sum_{i=1}^{M}\lambda_{i}u_{i}u_{i}^{*}$, where we let $M=2^{N}$. Observables are meant to correspond to physical properties that one can measure. The final ingredient of the model is Born’s rule, which associates a state-observable pair to a random measurement. More precisely, Born’s rule stipulates that measuring the physical property described by $A$ in the state described by |$\outerproduct{\psi}{\psi}$ | is equivalent to drawing from the random variable $X$ supported on the spectrum $\Lambda=\{\lambda_{1},\dots,\lambda_{M}\}\subset\mathbb{R}$ of $A$, and described by
$\mathbb{E}h(X)=\mathrm{Tr}\left(\outerproduct{\psi}{\psi}\sum_{i=1}^{M}h(\lambda_{i})u_{i}u_{i}^{*}\right)=\sum_{i=1}^{M}h(\lambda_{i})|\innerproduct{\psi}{u_{i}}|^{2},$ (11.1)
where $h$ is any real-valued function on $\Lambda$. Furthermore, independently preparing |$\ket{\psi}$ and measuring $A$ several times in a row yields statistically independent draws of the random variable $X$.
@@ 11.2 The canonical anticommutation relations lead to DPPs
In the rest of this section, we describe a state |$\ket{\psi}$ and an observable $A$ on $\mathbb{H}=(\mathbb{C}^{2})^{\otimes N}$ that can be efficiently prepared and measured on a quantum computer, and such that $X$ in (11.1) is a user-defined projection DPP on $\{1,\dots,N\}$. Generic DPPs on finite ground sets with Hermitian kernels, and more generally Pfaffian point processes, can also be obtained [BFF24], but we focus here on projection DPPs for simplicity.
@@
where $\{u,v\}\triangleq uv+vu$ is the anti-commutator of operators $u$ and $v$. Assuming the existence of a set of operators satisfying (11.2) for the moment, one can show [Nie05] the existence of a simultaneous eigenvector to all $c_{i}^{*}c_{i}$, which we denote by $\ket{\emptyset}$, always with eigenvalue zero. Moreover, the vectors
|$\ket{\mathbf{n}}=\prod_{i=1}^{N}(c_{i}^{*})^{n_{i}}\ket{\emptyset},\quad\mathbf{n}\in\{0,1\}^{N},$
@@
Coming back to the existence of operators satisfying (11.2), an example is given by the Jordan-Wigner operators $a_{1},\dots,a_{N}$; see e.g. Definition 3.3 in [BFF24]. The Jordan-Wigner operators come with additional tractability, in the sense that any product of $a_{i}$s applied to $\ket{\emptyset}$ is easy to prepare on a quantum computer, and any observable of the form $a_{i}^{*}a_{i}$ is easy to measure in such a state.
@@
One can show that $b_{1},\dots,b_{k}$ also satisfy (11.2). Now, consider the state |$\outerproduct{\psi}{\psi}$ |, where
|$\ket{\psi}=b_{1}^{*}\dots b_{r}^{*}\ket{\emptyset},$
and consider the observable $A=\prod_{i=1}^{N}a_{i}^{*}a_{i}$. Since all $a_{i}^{*}a_{i}$ are simultaneously diagonalizable in the Fock basis, they commute. In particular, $A$ is self-adjoint, making it a bona fide observable. Decomposing the trace appearing in Born’s rule (11.1) in the Fock basis, and carefully applying the canonical anti-commutation relations, one obtains a random variable $X$ that is a projection DPP with kernel $K=V_{[r],:}^{*}V_{[r],:}$. The determinantal character of the distribution precisely comes from this manipulation of anticommuting operators in Born’s rule, which physicists have come to call Wick’s theorem. We refer to Appendix A.7 of [BFF24] for a detailed computation, with the same notations as we used here.
Figure 4: A circuit sampling a DPP with projection kernel of rank $r=3$, with $N=5$ items. On the left-hand side, Pauli $X$ gates are used to create fermionic modes on the three first qubits. Note also the parallel QR Givens rotations on neighbouring qubits indicated by parametrized $XX+YY$ gates. On the right-hand side, measurements of occupation numbers are denoted by black squares (c.f. [BFF24]).
In addition to a more extensive introduction to the quantum computing tools used here, the paper [BFF24] also contains a derivation of state-observable pairs corresponding to generic DPPs with Hermitian kernels and Pfaffian point processes, as well as experimental demonstrations on IBM quantum computers.
@@ 12 Acknowledgements
SG was supported in part by the MOE grants R-146-000-250-133, R146-000-312-114, A-8002014-00-00 and MOE-T2EP20121-0013. H.S.T. was supported by the NUS Research Scholarship. VP and RB were supported by grants ERC-2019-STG-851866 and ANR20-CHIA-0002. The authors would like to thank Kin Aun Tan and Clarence Chew from the National University of Singapore for their comments and feedback on the paper.

### v4

@@ 11.1 The formalism of quantum physics
For simplicity, we henceforth restrict to the $2^{N}$-dimensional example $\mathbb{H}=(\mathbb{C}^{2})^{\otimes N}$, for some $N\geq 0$. This is the Hilbert space that is used in quantum computing to describe a system of $N$ qubits, i.e. $N$ particles that can each assume two states. In this finite-dimensional setting, an observable $A$ is simply a self-adjoint matrix, so that it diagonalizes in an orthonormal basis $A=\sum_{i=1}^{M}\lambda_{i}u_{i}u_{i}^{*}$, where we let $M=2^{N}$. Observables are meant to correspond to physical properties that one can measure. The final ingredient of the model is Born’s rule, which associates a state-observable pair to a random measurement. More precisely, Born’s rule stipulates that measuring the physical property described by $A$ in the state described by $|\outerproduct{\psi}{\psi}$ is equivalent to drawing from the random variable $X$ supported on the spectrum $\Lambda=\{\lambda_{1},\dots,\lambda_{M}\}\subset\mathbb{R}$ of $A$, and described by
[…]
@@ 11.2 The canonical anticommutation relations lead to DPPs
[…]
@@
[…]
@@
[…]
@@
One can show that $b_{1},\dots,b_{k}$ also satisfy (11.2). Now, consider the state $|\outerproduct{\psi}{\psi}$, where
$|,\ket{\psi}=b_{1}^{*}\dots b_{r}^{*}\ket{\emptyset}$,
and consider the observable $A=\prod_{i=1}^{N}a_{i}^{*}a_{i}$. Since all $a_{i}^{*}a_{i}$ are simultaneously diagonalizable in the Fock basis, they commute. In particular, $A$ is self-adjoint, making it a bona fide observable. Decomposing the trace appearing in Born’s rule (11.1) in the Fock basis, and carefully applying the canonical anti-commutation relations, one obtains a random variable $X$ that is a projection DPP with kernel $K=V_{[r],:}^{*}V_{[r],:}$. The determinantal character of the distribution precisely comes from this manipulation of anticommuting operators in Born’s rule, which physicists have come to call Wick’s theorem. We refer to Appendix A.7 of [BaFaFe24] for a detailed computation, with the same notations as we used here.
Figure 4: A circuit sampling a DPP with projection kernel of rank $r=3$, with $N=5$ items. On the left-hand side, Pauli $X$ gates are used to create fermionic modes on the three first qubits. Note also the parallel QR Givens rotations on neighbouring qubits indicated by parametrized $XX+YY$ gates. On the right-hand side, measurements of occupation numbers are denoted by black squares (c.f. [BaFaFe24]).
In addition to a more extensive introduction to the quantum computing tools used here, the paper [BaFaFe24] also contains a derivation of state-observable pairs corresponding to generic DPPs with Hermitian kernels and Pfaffian point processes, as well as experimental demonstrations on IBM quantum computers.
@@ 12 Acknowledgements
SG was supported in part by the MOE grants R-146-000-250-133, R146-000-312-114, A-8002014-00-00 and MOE-T2EP20121-0013. H.S.T. was supported by the National University of Singapore and MOE of Singapore through the grant A-8003576-00-00. VP and RB were supported by grants ERC-2019-STG-851866 and ANR20-CHIA-0002. The authors would like to thank Kin Aun Tan and Clarence Chew from the National University of Singapore for their comments and feedback on the paper.